# Dynamic Pricing with Adversarially-Censored Demands

### v1

@@ 7 Conclusions
In this paper, we studied the online pricing problem with adversarial inventory constraints imposed over time series. We introduced an optimistic strategy and a C20CB algorithm that is capable of approaching the optimal prices from inventory-censored demands. Our algorithm enjoys a regret guarantee at $\tilde{O}(\sqrt{T})$ with high probability, which is information-theoretically optimal. To the best of our knowledge, we are the first to address this adversarial-inventory pricing problem, and our results indicate that the demand-censoring effect does not substantially increase the hardness of pricing in terms of minimax regret.
Appendix A Proof Details
A.1 Proof of Lemma 5.2
Proof.
@@ Appendix A Proof Details
[…]
@@
    $=p\gamma_{t}-p(\gamma_{t}-a+bp)F(\gamma_{t}-a+bp)-p(c-G(c)-F(\gamma_{t}-a+bp)\cdot(\gamma_{t}-a+bp)+G(\gamma_{t}-a+bp))$
@@
[…]
@@
[…]
@@
    $r^{\prime}_{t}(p)\equiv 0,\forall p\in[p_{t}^{*},q_{t}^{*}]\text{ due to the monotonicity of }r^{\prime}_{t}(p)$ (17)
@@
    $\Rightarrow$ $0\displaystyle 0\leq F(\gamma_{t}-a+bp_{t}^{*})\leq\lim_{p\rightarrow q_{t}^{*}-}F(\gamma_{t}-a+bp)=$ 0
@@
[…]
@@
    Let $C_{\epsilon}:=\frac{b_{\min}}{2}\cdot\inf_{\gamma_{t}\in[\gamma_{\min},\gamma_{\max}]}F(\gamma_{t}-a+bp_{t}^{*})$. As $[\gamma_{\min},\gamma_{\max}]$ is a close set and $F(\gamma_{t}-a+bp_{t}^{*})$ holds for any $\gamma_{t}\in[\gamma_{\min},\gamma_{\max}]$, we know that $C_{\epsilon}>0$ is a universal constant. Given this coefficient, for any $p_{1},p_{2}\in[p_{t}^{*}-\epsilon_{t},p_{t}^{*}+\epsilon_{t}],p_{1}<p_{2}$, we have
@@
A.2 Proof of Lemma 5.3
Proof.
@@
Here the first row is due to Law of Total Expectation, and the last row is due to the zero-mean assumption of $N_{t}$ (see Assumption 3.6). Given this equation, we have:
@@
Here the fifth row requires $\frac{1}{8bp_{\max}}\geq\frac{1}{\gamma_{\min}-\gamma_{0}}\sqrt{\frac{1}{2}\log\frac{2}{\eta\delta}}\cdot\frac{1}{\sqrt{\tau}}$, which further requires $T\geq(\frac{8bp_{\max}}{\gamma_{\min}-\gamma_{0}})^{4}\cdot\frac{1}{4}\log^{2}\frac{2}{\eta\theta}$. According to Assumption 3.8, we know that this inequality holds. Denote $C_{b}:=\frac{4b_{\max}^{2}p_{\max}}{\gamma_{\min}-\gamma_{0}}\cdot\sqrt{\frac{1}{2}\log\frac{2}{\eta\delta}}$ and we have $|\hat{b}-b|\leq C_{b}\cdot\frac{1}{\sqrt{\tau}}$ with high probability.
@@
Denote $C_{a}:=p_{\max}(C_{b}+b_{\max}\sqrt{\frac{1}{2}\log\frac{2}{\eta\delta}})$ and the lemma is proven. ∎
A.3 Proof of Lemma 5.4
Proof.
Here we consider the time periods before time $t$, and we use an index $s$ to denote each time period $s=1,2,\ldots,t-1,t$. As a consequence, we have the notations $D_{s},\gamma_{s},k_{s},\mathds{1}_{s}$ corresponding to $D_{t},\gamma_{t},k_{t},\mathds{1}_{t}$ as we defined in Section 3 and Section 4. Also, we denote $N_{k}(t),F_{k}(t)$ and $G_{k}(t)$ as the value of $N_{k},F_{k}$ and $G_{k}$ at the beginning of time period $t$.
From Algorithm 1, we have
@@
$=$ $\frac{1}{N_{k}(t)}\sum_{s=1}^{t-1}\mathds{1}[k_{s}==k](G(c)-G(\gamma_{s}-a+b\cdot\frac{2k\Delta-\gamma_{s}+\hat{a}}{\hat{b}})).$
@@
$|G(\gamma_{s}-a+b\cdot\frac{2k\Delta-\gamma_{s}+\hat{a}}{\hat{b}})|$ (32)
$=$ $|G(\gamma_{s}-a+b\cdot\frac{2k\Delta-\gamma_{s}+\hat{a}}{b}+b(\frac{1}{\hat{b}}-\frac{1}{b})(k\delta-\gamma_{s}+\hat{a}))-G(2k\Delta)|$
@@
$|\mathbb{E}[G_{k}(t)]-(G(c)-G(2k\Delta))|\leq$ $\frac{1}{N_{k}(t)}\sum_{s=1}^{t-1}\mathds{1}[k_{s}==k](C_{a}+C_{b}\cdot\frac{c+a_{\max}}{b_{\min}})\cdot\frac{1}{\sqrt{\tau}}$ (33)
@@
$=$ $\frac{1}{N_{k}(t)}\sum_{s=1}^{t-1}\mathds{1}[k_{s}==k]\cdot F(\gamma_{s}-a+b\cdot\frac{2k\Delta-\gamma_{s}+\hat{a}}{\hat{b}}).$
Similar to Eq. 32, since $F(x)$ is $L_{F}$-Lipschitz, we have
@@
$+|G(2k\Delta)-G(\gamma_{t}-a+b\cdot\frac{2k\Delta-(\gamma_{t}-\hat{a})}{\hat{b}})|+bp_{k,t}|F(2k\Delta)-F(\gamma_{t}-a+b\cdot\frac{2k\Delta-(\gamma_{t}-\hat{a})}{\hat{b}})|$
@@
$+|G(2k\Delta)-G(\gamma_{t}-a+b\cdot\frac{2k\Delta-(\gamma_{t}-\hat{a})}{\hat{b}})|+b_{\max}p_{\max}\cdot|F(2k\Delta)-F(\gamma_{t}-a+b\cdot\frac{2k\Delta-(\gamma_{t}-\hat{a})}{\hat{b}})|$
@@
=:absent:=:= : $C_{N}\cdot\frac{1}{\sqrt{N_{k}(t)}}+C_{\tau}\cdot\frac{1}{\sqrt{\tau}}=\Delta_{k}.$
@@
$C_{N}$ $:=(c+a_{\max}+b_{\max}p_{\max})\cdot\sqrt{\frac{1}{2}\log\frac{2}{\eta\delta}},$ (40)
$C_{\tau}$ $:=2(b_{\max}p_{\max}L_{F}+1)(C_{a}+C_{b}\cdot\frac{c+a_{\max}}{b_{\min}})+p_{\max}C_{b}.$
Finally, we apply the union bound on the probability, and know that Eq. 39 holds with probability $\Pr\geq 1-6\eta\delta$. ∎
A.4 Proof of Lemma 5.5
Proof.
We first prove the lemma under Case 1 when some confidence bound contains 00. Denote $\rho_{t}:=\min\{|r^{\prime}_{t}(p_{t}^{*}-\epsilon_{t})|,|r^{\prime}_{t}(p_{t}^{*}+\epsilon_{t})|\}$, and we know that $\rho_{t}>0$ due to the uniqueness of $p_{t}^{*}$.
Now, let $N_{0}:=\frac{36C_{N}^{2}}{\rho_{t}^{2}}$, where $C_{N}$ is the constant coefficient define in Lemma 5.4. Given this, when $N_{k}(t)\geq N_{0}$, we have
@@
Also, since $T$ is assumed as larger than any constant (see Assumption 3.8), we have $T\geq\frac{1296C_{\tau}^{4}}{\rho_{t}^{4}}$, and therefore
@@
Given eq. 41 and eq. 42, we know that $\Delta_{k}(t)=C_{N}\cdot\frac{1}{\sqrt{N_{k}(t)}}+C_{\tau}\cdot\frac{1}{\sqrt{\tau}}\leq\frac{\rho_{t}}{6}+\frac{\rho_{t}}{6}=\frac{\rho_{t}}{3}$. Now, if $0\in[\hat{r}_{k,t}-\Delta_{k}(t),\hat{r}_{k,t}+\Delta_{k}(t)]$, we have $|\hat{r}_{k,t}|\leq\Delta_{k}(t)$ and therefore
@@
Since $r^{\prime}_{t}(p)$ is monotonically non-increasing, any $p\in[0,p_{\max}]$ satisfying $r^{\prime}_{t}(p)<\rho_{t}$ should satisfy $p\in(p_{t}^{*}-\epsilon_{t},p_{t}^{*}+\epsilon_{t})$. Therefore, we have $p_{k,t}\in(p_{t}^{*}-\epsilon_{t},p_{t}^{*}+\epsilon_{t})$. According to Lemma 5.2 Property (7), we have:
@@
Let $C_{in}:=8C_{v}\max\{C_{N}^{2},C_{\tau}^{2}\}$ and the first part of Lemma 5.5 holds.
@@
Denote $\theta_{t}:=\inf_{k}\min\{|\hat{r}_{k,t}+\Delta_{k}(t)|,|\hat{r}_{k,t}-\Delta_{k}(t)|\}$, and we know that $\min\{|\hat{r}_{k_{t},t}+\Delta_{k}(t)|,|\hat{r}_{k_{t},t}-\Delta_{k}(t)|\}\theta_{t}>0$, where $k_{t}$ is the $k$ such that $p_{k,t}$ is proposed at time $t$. Therefore, we have $|\hat{r}_{k,t}+\Delta_{k}(t)|\geq\theta_{t}$ and $|\hat{r}_{k,t}-\Delta_{k}(t)|\geq\theta_{t},\forall k$. According to the prerequisite of Lemma 5.5 Part (2), there exists $k_{0}$ such that
[…]
@@
As $T$ is sufficiently large, we have $\frac{L_{r}}{b_{\min}}(C_{a}+C_{b}p_{\max})\cdot\frac{1}{T^{1/4}}\leq\frac{1}{6}\cdot\rho_{t}$ where $\rho_{t}:=\min\{|r^{\prime}_{t}(p_{t}^{*}-\epsilon_{t})|,|r^{\prime}_{t}(p_{t}^{*}+\epsilon_{t})|\}$. Let $N_{1}:=\frac{144C_{N}^{2}}{\rho_{t}^{2}}$. Similar to Eq. 41 and Eq. 42, we have
@@
Since $r^{\prime}_{t}(p)$ is monotonically non-increasing, we know that $p_{k_{t},t}\in(p_{t}^{*}-\epsilon_{t},p_{t}^{*}+\epsilon_{t})$ similar to the analysis in Case (1), and again we have $|r_{t}(p_{t}^{*})-r_{t}(p_{k_{t},t})|\leq 8C_{v}(C_{N}^{2}\cdot\frac{1}{N_{k_{t}}(t)}+C_{\tau}^{2}\cdot\frac{1}{\tau})$. Now we complete the proof of Lemma 5.5 on both circumstances. ∎
A.5 Proof of Lemma 5.6
Proof.
When $\gamma_{t}>\frac{\hat{a}+C_{a}\cdot\frac{1}{\sqrt{\tau}}}{2}+c>\frac{a}{2}+c$, we know that the demand at $p=\frac{a}{2b}$ and its neighborhood is not censored. As a result, $\frac{a}{2b}$ is still a local optimal (and therefore global optimal due to the unimodality) of $r_{t}(p)$. According to Lemma 5.3, we have:
$r_{t}(\frac{a}{2b})-r_{t}(\frac{\hat{a}}{2\hat{b}})$ $\leq C_{s}(\frac{a}{2b}-\frac{\hat{a}}{2\hat{b}})^{2}$ (50)

### v2

@@ 7 Conclusions
In this paper, we studied the online pricing problem with adversarial inventory constraints imposed over time series. We introduced an optimistic strategy and a C20CB algorithm that is capable of approaching the optimal prices from inventory-censored demands. Our algorithm achieves a regret guarantee of $\tilde{O}(\sqrt{T})$ with high probability, which is information-theoretically optimal. To the best of our knowledge, we are the first to address this adversarial-inventory pricing problem, and our results indicate that the demand-censoring effect does not substantially increase the hardness of pricing in terms of minimax regret.
Acknowledgement
Jianyu Xu started this work as a Ph.D. student at University of California, Santa Barbara. Xi Chen would like to thank the support from NSF via the Grant IIS-1845444.
Appendix
Appendix 0.A More Related Works
Here we discuss more related works as a complement to Section˜2.
Contextual pricing: Linear valuation and binary-censored demand.
A surge of research has focused on feature-based (or contextual) dynamic pricing [21, 2, 43, 39]. These works considered situations where each pricing period is preceded by a context, influencing both the demand curve and noise distribution. Specifically, [21, 31, 61] explored a linear valuation framework with known distribution noise, leading to binary customer demand outcomes based on price comparisons to their valuations. Expanding on this, [29, 26, 41, 62] examined similar models but with unknown noise distributions. In another vein, [6, 54, 63] investigated personalized pricing where demand is modeled as a generalized linear function sensitive to contextual price elasticity. Many of these works on valuation-based contextual pricing also assumed a censored demand: The seller only observes a binary feedback determined by a comparison of price with valuation, instead of observing the valuation directly. However, it is important to differentiate between the linear (potential) demand model we assumed and their linear valuation models, and there exists no inclusive relationship to each other.
Dynamic pricing under constraints.
A variety of research works have been devoted to the field of dynamic pricing under specific concerns, which restricts the stages and outcomes of the market. This includes resource allocation [16, 53, 60] and price discrimination [18, 22, 20, 59, 32] as two typical instances. In this work, the constraints are twofold: First, it blocks our observations to the real potential demand, leading to biased estimates. Second, it restricts us from fulfilling the potential demand, leading to shifted targets (the optimal prices).
Multi-armed bandits.
Multi-armed bandits (MAB) formalize sequential decision-making under uncertainty via the exploration–exploitation tradeoff, originating in the stochastic setting studied by [37]. Modern finite-time analyses led to widely used algorithms such as UCB [3], Thompson Sampling [51], and the EXP-3/4 family [4]. Contextual bandits extend MAB by conditioning rewards on observed features, enabling personalization and decision-making with side information. Representative algorithmic and theoretical foundations include Epoch greedy [38], LinUCB [19] and Taming-the-monster [1]. [5] introduces a “Bandits with knapsacks” model that incorporates resource and budget limitations. A separated stream of “Continuum-armed bandits”, introduced by [35], generalize discrete action sets to continuous domains and exploit smoothness/metric structure. Their method is applicable to our problem setting but will lead to a sub-optimal $O(T^{3/5})$ regret. A recent model of “Generative bandits” [58] allows new arms to be generated on the fly (at a certain cost) for free future reuse. From a more general perspective, the stream of works in “Zeroth-order optimization (ZOO)” [56] also belongs to bandits optimization.
Appendix 0.B Proof Details
0.B.1 Proof of Lemma˜1
Proof
@@ 0.B.1 Proof of Lemma˜1
    $=p\left(\int_{-c}^{\gamma_{t}-a+bp}(a-bp+x)f(x)dx+\int_{\gamma_{t}-a+bp}^{c}\gamma_{t}f(x)dx\right)$
    $=p\left(\int_{-c}^{c}(a-bp+x)f(x)dx+\int_{\gamma_{t}-a+bp}^{c}(\gamma_{t}-(a-bp+x))f(x)dx\right)$
    $=p\left((a-bp)\cdot\int_{-c}^{c}f(x)dx+\int_{-c}^{c}xf(x)dx\right$.
    $\quad\left.+(\gamma_{t}-a+bp)\int_{\gamma_{t}-a+bp}^{c}f(x)dx-\int_{\gamma_{t}-a+bp}^{c}xf(x)dx\right)$
@@
    $=p\gamma_{t}-p(\gamma_{t}-a+bp)F(\gamma_{t}-a+bp)$
    $\quad-p(c-G(c)-F(\gamma_{t}-a+bp)\cdot(\gamma_{t}-a+bp)+G(\gamma_{t}-a+bp))$
@@
    Here we adopt the notation $f(x)$ as proximal derivatives of $F(x)$. According to Rademacher’s Theorem (see Section 3.5 of [27]), given that $F(x)$ is Lipschitz, the measure of $x$ such that $f(x)$ does not exist is zero, hence the integral holds. Here the eighth line comes from $\int_{-c}^{c}f(x)dx=F(c)-F(-c)=1$ and $\int_{-c}^{c}xf(x)dx=\mathbb{E}[x]=0$. Given the close form of $r_{t}(p)$, we derive the form of $r^{\prime}_{t}(p)$.
@@
    We first show the existence of $p_{t}^{*}\in[0,\frac{a}{b}]$ such that $r^{\prime}_{t}(p_{t}^{*})=0$. Recall that $G(c+x)=G(c)+x$ for $\forall x>0$, and $G(c)-G(-c)=\int_{-c}^{c}F(\omega)d\omega\geq 0$, and that $\gamma_{t}>2c>c$ as we assumed. Given those, we have:
@@
[…]
@@
    $\Rightarrow$ $0\leq F(\gamma_{t}-a+bp_{t}^{*})\leq\lim_{p\rightarrow q_{t}^{*}-}F(\gamma_{t}-a+bp)=0$
@@
    $\qquad\quad=\gamma_{t}-c+G(c)-\int_{-\infty}^{\gamma_{t}-a+bp_{t}^{*}}F(\omega)d\omega-0$
@@
    Let $C_{\epsilon}:=\frac{b_{\min}}{2}\cdot\inf_{\gamma_{t}\in[\gamma_{\min},\gamma_{\max}]}F(\gamma_{t}-a+bp_{t}^{*})$. As $\gamma_{\min},\gamma_{\max}]$ is a close set and $F(\gamma_{t}-a+bp_{t}^{*})$ holds for any $\gamma_{t}\in[\gamma_{\min},\gamma_{\max}]$, we know that $C_{\epsilon}>0$ is a universal constant. Given this coefficient, for any $p_{1},p_{2}\in[p_{t}^{*}-\epsilon_{t},p_{t}^{*}+\epsilon_{t}],p_{1}<p_{2}$, we have
@@
0.B.2 Proof of Lemma˜2
Proof
@@ 0.B.2 Proof of Lemma˜2
Here the first row is due to Law of Total Expectation, and the last row is due to the zero-mean assumption of $N_{t}$ (see Assumption˜3). Given this equation, we have:
@@
Here the fifth row requires $\frac{1}{8bp_{\max}}\geq\frac{1}{\gamma_{\min}-\gamma_{0}}\sqrt{\frac{1}{2}\log\frac{2}{\eta\delta}}\cdot\frac{1}{\sqrt{\tau}}$, which further requires $T\geq(\frac{8bp_{\max}}{\gamma_{\min}-\gamma_{0}})^{4}\cdot\frac{1}{4}\log^{2}\frac{2}{\eta\theta}$. According to Assumption˜5, we know that this inequality holds. Denote $C_{b}:=\frac{4b_{\max}^{2}p_{\max}}{\gamma_{\min}-\gamma_{0}}\cdot\sqrt{\frac{1}{2}\log\frac{2}{\eta\delta}}$ and we have $|\hat{b}-b|\leq C_{b}\cdot\frac{1}{\sqrt{\tau}}$ with high probability.
@@
Denote $C_{a}:=p_{\max}(C_{b}+b_{\max}\sqrt{\frac{1}{2}\log\frac{2}{\eta\delta}})$ and the lemma is proven.
0.B.3 Proof of Lemma˜3
Proof
Here we consider the time periods before time $t$, and we use an index $s$ to denote each time period $s=1,2,\ldots,t-1,t$. As a consequence, we have the notations $D_{s},\gamma_{s},k_{s},\mathds{1}_{s}$ corresponding to $D_{t},\gamma_{t},k_{t},\mathds{1}_{t}$ as we defined in Section˜3 and Section˜4. Also, we denote $N_{k}(t),F_{k}(t)$ and $G_{k}(t)$ as the value of $N_{k},F_{k}$ and $G_{k}$ at the beginning of time period $t$.
From Algorithm˜1, we have
@@ 0.B.3 Proof of Lemma˜3
$=$ $).\displaystyle\frac{1}{N_{k}(t)}\sum_{s=1}^{t-1}\mathds{1}[k_{s}==k](G(c)-G(\gamma_{s}-a+b\cdot\frac{2k\Delta-\gamma_{s}+\hat{a}}{\hat{b}}$)).
@@
$|G(\gamma_{s}-a+b\cdot\frac{2k\Delta-\gamma_{s}+\hat{a}}{\hat{b}})-G(2k\Delta)|$ (36)
$=$ $|G(\gamma_{s}-a+b\cdot\frac{2k\Delta-\gamma_{s}+\hat{a}}{b}+b(\frac{1}{\hat{b}}-\frac{1}{b})(2k\delta-\gamma_{s}+\hat{a}))-G(2k\Delta)|$
@@
$|\mathbb{E}[G_{k}(t)]-(G(c)-G(2k\Delta))|\leq$ $\frac{1}{N_{k}(t)}\sum_{s=1}^{t-1}\mathds{1}[k_{s}==k]|G(\gamma_{s}-a+b\cdot\frac{2k\Delta-\gamma_{s}+\hat{a}}{\hat{b}})-G(2k\Delta)|$ (37)
$\leq$ $\frac{1}{N_{k}(t)}\sum_{s=1}^{t-1}\mathds{1}[k_{s}==k](C_{a}+C_{b}\cdot\frac{c+a_{\max}}{b_{\min}})\cdot\frac{1}{\sqrt{\tau}}$
@@
[…]
Similar to Eq.˜36, since $F(x)$ is $L_{F}$-Lipschitz, we have
@@
[…]
@@
[…]
@@
$=:$ $C_{N}\cdot\frac{1}{\sqrt{N_{k}(t)}}+C_{\tau}\cdot\frac{1}{\sqrt{\tau}}=\Delta_{k}.$
@@
$C_{N}$ $=(c+a_{\max}+b_{\max}p_{\max})\cdot\sqrt{\frac{1}{2}\log\frac{2}{\eta\delta}},$ (44)
$C_{\tau}$ $=2(b_{\max}p_{\max}L_{F}+1)(C_{a}+C_{b}\cdot\frac{c+a_{\max}}{b_{\min}})+p_{\max}C_{b}.$
Finally, we apply the union bound on the probability, and know that Eq.˜43 holds with probability $\Pr\geq 1-6\eta\delta$.
0.B.4 Proof of Lemma˜4
Proof
We first prove the lemma under Case 1 when some confidence bound contains 0. Denote $\rho_{t}:=\min\{|r^{\prime}_{t}(p_{t}^{*}-\epsilon_{t})|,|r^{\prime}_{t}(p_{t}^{*}+\epsilon_{t})|\}$, and we know that $\rho_{t}>0$ due to the uniqueness of $p_{t}^{*}$.
Now, let $N_{0}:=\frac{36C_{N}^{2}}{\rho_{t}^{2}}$, where $C_{N}$ is the constant coefficient define in Lemma˜3. Given this, when $N_{k}(t)\geq N_{0}$, we have
@@ 0.B.4 Proof of Lemma˜4
Also, since $T$ is assumed as larger than any constant (see Assumption˜5), we have $T\geq\frac{1296C_{\tau}^{4}}{\rho_{t}^{4}}$, and therefore
@@
Given eq.˜45 and eq.˜46, we know that $\Delta_{k}(t)=C_{N}\cdot\frac{1}{\sqrt{N_{k}(t)}}+C_{\tau}\cdot\frac{1}{\sqrt{\tau}}\leq\frac{\rho_{t}}{6}+\frac{\rho_{t}}{6}=\frac{\rho_{t}}{3}$. Now, if $0\in[\hat{r}_{k,t}-\Delta_{k}(t),\hat{r}_{k,t}+\Delta_{k}(t)]$, we have $|\hat{r}_{k,t}|\leq\Delta_{k}(t)$ and therefore
@@
Since $r^{\prime}_{t}(p)$ is monotonically non-increasing, any $p\in[0,p_{\max}]$ satisfying $r^{\prime}_{t}(p)<\rho_{t}$ should satisfy $p\in(p_{t}^{*}-\epsilon_{t},p_{t}^{*}+\epsilon_{t})$. Therefore, we have $p_{k,t}\in(p_{t}^{*}-\epsilon_{t},p_{t}^{*}+\epsilon_{t})$. According to Lemma˜1 Property (7), we have:
@@
Let $C_{in}:=8C_{v}\max\{C_{N}^{2},C_{\tau}^{2}\}$ and the first part of Lemma˜4 holds.
@@
Denote $\theta_{t}:=\inf_{k}\min\{|\hat{r}_{k,t}+\Delta_{k}(t)|,|\hat{r}_{k,t}-\Delta_{k}(t)|\}$, and we know that $\min\{|\hat{r}_{k_{t},t}+\Delta_{k}(t)|,|\hat{r}_{k_{t},t}-\Delta_{k}(t)|\}\theta_{t}>0$, where $k_{t}$ is the $k$ such that $p_{k,t}$ is proposed at time $t$. Therefore, we have $|\hat{r}_{k,t}+\Delta_{k}(t)|\geq\theta_{t}$ and $|\hat{r}_{k,t}-\Delta_{k}(t)|\geq\theta_{t},\forall k$. According to the prerequisite of Lemma˜4 Part (2), there exists $k_{0}$ such that
$\left\{\begin{aligned} &\hat{r}_{k_{0},t}-\Delta_{k_{0}}(t)\geq\theta_{t}\\ &\hat{r}_{k_{0}+1,t}+\Delta_{k_{0}+1}(t)\leq\theta_{t}.\end{aligned}\right$. (50)
@@
As $T$ is sufficiently large, we have $\frac{L_{r}}{b_{\min}}(C_{a}+C_{b}p_{\max})\cdot\frac{1}{T^{1/4}}\leq\frac{1}{6}\cdot\rho_{t}$ where $\rho_{t}:=\min\{|r^{\prime}_{t}(p_{t}^{*}-\epsilon_{t})|,|r^{\prime}_{t}(p_{t}^{*}+\epsilon_{t})|\}$. Let $N_{1}:=\frac{144C_{N}^{2}}{\rho_{t}^{2}}$. Similar to Eq.˜45 and Eq.˜46, we have
@@
Since $r^{\prime}_{t}(p)$ is monotonically non-increasing, we know that $p_{k_{t},t}\in(p_{t}^{*}-\epsilon_{t},p_{t}^{*}+\epsilon_{t})$ similar to the analysis in Case (1), and again we have $|r_{t}(p_{t}^{*})-r_{t}(p_{k_{t},t})|\leq 8C_{v}(C_{N}^{2}\cdot\frac{1}{N_{k_{t}}(t)}+C_{\tau}^{2}\cdot\frac{1}{\tau})$. This completes the proof of Lemma˜4 on both circumstances.
0.B.5 Proof of Lemma˜5
Proof
When $\gamma_{t}>\frac{\hat{a}+C_{a}\cdot\frac{1}{\sqrt{\tau}}}{2}+c>\frac{a}{2}+c$, we know that the demand at $p=\frac{a}{2b}$ and its neighborhood is not censored. As a result, $\frac{a}{2b}$ is still a local optimal (and therefore global optimal due to the unimodality) of $r_{t}(p)$. According to Lemma˜2, we have:
[…]